There are 63 students in a middle school chorus. There are 11 more boys than girls. How many boys x and girls y are in the chorus?
- Thanks to anyone who answers.

Answer by stanbon(75887)

There are 63 students in a middle school chorus. There are 11 more boys than girls. How many boys x and girls y are in the chorus?
---------------------------------
b + g = 65
b = g + 11
------------
Substitute for "b" and solve for "g":
g+11+g = 65
2g = 54
g = 27 (# of girls)
---
Solve for "b":
b + 27 = 65
b = 38 (# of boys)
======================
Cheers,
Stan H.
======================